OpenStudy (anonymous):

Let f(A)= (cosA/1+sinA) + (1+sinA/cosA) Prove that f(A)=2secA

OpenStudy (koikkara):

@sianoroche = (2 + 2sinA) / (1+sinA)(cosA) = (2 + 2sinA)(secA) / (1+sinA) .....[1/cosA=secA] = 2(1+sinA)(secA) / (1+sinA) .....[Factor out the 2] = 2secA

OpenStudy (clalgee):

Welcome to Openstudy!!